OpenStudy (anonymous):

Consider the following pair of equations: y = 3x + 3 y = x − 1 Explain how you will solve the pair of equations by substitution. Show all the steps and write the solution in (x, y) form.

OpenStudy (anonymous):

3x + 3 = x - 1

OpenStudy (kidrah69):

yep go on

OpenStudy (kidrah69):

hint:subtract 3 from both sides

OpenStudy (anonymous):

kidrah69 your picture looks freaky

OpenStudy (kidrah69):

thanks @cami1231 did u do the step?

OpenStudy (anonymous):

3x = x - 2?

OpenStudy (kidrah69):

yep now subtract 1x from both sides

OpenStudy (anonymous):

2x = -2?

OpenStudy (kidrah69):

yep now divide 2x/-2

OpenStudy (anonymous):

-1x

OpenStudy (kidrah69):

thats sort of correct....you get X=-1 because when you divde it cancels out

OpenStudy (anonymous):

oo

OpenStudy (kidrah69):

\[y=3(-1)+3\] now solve this

OpenStudy (kidrah69):

to get the Y value

OpenStudy (kidrah69):

i plugged in what we got for the other equation in place of X

OpenStudy (anonymous):

y = -3 + 3

OpenStudy (kidrah69):

yep and what is -3+3?

OpenStudy (anonymous):

y = 0

OpenStudy (kidrah69):

yep so whats the final answer?

OpenStudy (anonymous):

(-1,0)